# Splitting obstructions and $\mathbb{Z}_{2}$ invariants in time-reversal symmetric topological insulators

### v1

@@ Appendix A Tools from algebraic geometry
This argument is already present in several works, for example in [graf2013bulk]. We prove it below for the sake of completeness.
@@
In fact, by definition, the two spans intersect only at 0 and if, by contradiction, there is a vector $v$ that cannot be obtained by adding elements of the two spans. Then, up to renormalization, the vector
@@
$\begin{split}T\tilde{v}&=Tv-\sum_{j\in J}\overline{\left\langle v_{j},v\right\rangle Tv_{j}}+\sum_{j\in J}\overline{\left\langle Tv_{j},v\right\rangle}v_{j}=\\ &=Tv-\sum_{j\in J}\left\langle Tv_{j},Tv\right\rangle Tv_{j}-\sum_{j\in J}\left\langle v_{j},Tv\right\rangle v_{j}\end{split}$
@@
Given a map $f:U(1)\to U(1)$, the winding number, often denoted $[f]$, is an integer number exhibited by the natural isomorphism $\pi_{1}(U(1))\simeq\mathbb{Z}$. It can be computed by considering the natural covering $g:\mathbb{R}\to U(1)$ such that $g(t)=e^{i2\pi t}$. Then, the map $f$ can be lifted to a continuous map $\mu:[-\pi,\pi]\to\mathbb{R}$ such that $g(\mu(t))=f(\pi(t))$. Despite the numerous lifting options, the quantity $f]=\mu(\pi)-\mu(-\pi)\in\mathbb{Z}$ is well defined and depends only on the homotopy class of $f$.
@@
If $\lambda:\mathbb{T}^{1}\to U(1)$ is a continuous map with $\lambda(\pi)=\lambda(-\pi)=\lambda(0)=1$ and such that $\lambda(-k)=\overline{\lambda(k)}$, then $\lambda]\in 2\mathbb{Z}$ is even.
@@
$\lambda]=\frac{\mu(\pi)-\mu(-\pi)}{2\pi}=\frac{2\mu(\pi)}{2\pi}\in 2\mathbb{Z}$
@@
In fact, the two branches of the function connect for $k=(t-1)\pi,(1-t)\pi$ and also connect for $k=\pi=-\pi$ since $\lambda(\pi)=\lambda(-\pi)$. This means that $\lambda_{t}(k)$ is continuous. ∎
Corollary A.5.
Consider $U(1)=\left\{e^{i\theta}\in\mathbb{C}\right\}$ as a topological group with the standard product of $\mathbb{C}$. If $f,g:U(1)\to U(1)$ are two smooth maps with $f(1)=g(1)=1$ then the winding number of $f\cdot g$ is the sum of the winding number of $f$ and the one of $g$.
@@
If we apply the previous theorem we know that $f\cdot g$ is homotopic to fgf*g and it is well known that $\pi_{1}(U(1),1)$ is isomorphic as a group to $\mathbb{Z}$ with the group isomorphism that maps every function to its winding number (see e.g. [may1999concise, Chapter 1, Section 5]). So, if we denote by $[h]$ the winding number of the function $h$, we have that:
$f\cdot g]=[f*g]=[f]+[g].\qed$
Theorem A.6.
If $J=\begin{pmatrix}0&-\mathbb{I}_{n}\\ \mathbb{I}_{n}&0\end{pmatrix}$, $\mathcal{K}$ is the standard complex conjugation and $\mathrm{Sp}(2n)=\{U\in U(2n)|J\mathcal{K}U=UJ\mathcal{K}\}$, then $\mathrm{Sp}(2n)$ is simply connected.
Proof.
A matrix $U$ is in $\mathrm{Sp}(2n)$ if and only if it can be written in exponential form using $U=e^{iH}$, where $H$ is self-adjoint such that $(JH)^{t}=JH$ thanks to Proposition 3.5 and Corollary 11.10 in [hall2013lie]. Since this relation is linear, we can move $U$ toward the identity with the continuous transformation $U_{t}=e^{itH}$ for $t\in[0,1]$. ∎
Theorem A.7.
If $\mathcal{H}$ is a complex Hilbert space with finite dimension, then $\pi_{0}(\mathcal{U}(\mathcal{H}))\simeq\{0\}$, $\pi_{1}(\mathcal{U}(\mathcal{H}))\simeq\mathbb{Z}$ using the homomorphism that computes the winding number of the determinant and $\pi_{2}(\mathcal{U}(\mathcal{H}))$.
Proof.
We can express $\mathcal{U}(\mathcal{H})$ as $\mathrm{U}(n)$, set of $n\times n$ unitary and complex matrices. Then we can express it as a Cartesian product using the bijective map:
$\begin{matrix}\mathrm{U}(n)&\to&S^{1}\times\mathrm{SU}(n)\\ U&\mapsto&\left(\det(U),\begin{pmatrix}\det(U)^{-1}&0\\ 0&\mathbb{I}_{n-1}\end{pmatrix}U\right)\end{matrix}$
This means that there is a short exact sequence of groups
$\{\mathbb{I}\}\to\mathrm{SU}(n)\to\mathrm{U}(n)\to S^{1}\to\{\mathbb{I}\}.$
A sequence like this induces a long exact sequence
$\cdots\to\pi_{j+1}(S^{1})\to\pi_{j}(\mathrm{SU}(n))\to\pi_{j}(\mathrm{U}(n))\to\pi_{j}(S^{1})\to\pi_{j-1}(\mathrm{SU}(n))\to\cdots$
Since $S^{1}$ admits a universal covering using $\mathbb{R}$, we have that $\pi_{j}(S^{1})$ is trivial for $j<1$. Moreover, using the table containing the values of $\pi_{j}(\mathrm{SU}(n))$ in Appendix A, Section 6, Part VII of [ito1993encyclopedic] we have that:
$\pi_{0}(S^{1})\simeq\pi_{2}(S^{1})\simeq\pi_{0}(\mathrm{SU}(n))\simeq\pi_{1}(\mathrm{SU}(n))\simeq\pi_{2}(\mathrm{SU}(n))\simeq\{0\},$
and also that $\pi_{1}(S^{1})\simeq\mathbb{Z}$ using the homomorphism that computes the winding number of the map. Additional details on this topic can be found in [hatcher2005algebraic] on page 344. ∎
Theorem A.8 (Kuiper’s theorem).

### v2

@@ Appendix A Tools from algebraic geometry
This argument is already present in several works, for example in [gontier2022symmetric]. We prove it below for the sake of completeness.
@@
In fact, by definition, the two spans intersect only at 0 and if, by contradiction, there is a vector $v$ that cannot be obtained by adding elements of the two spans, then, up to renormalization, the vector
@@
$\begin{split}T\tilde{v}&=Tv-\sum_{j\in J}\overline{\left\langle v_{j},v\right\rangle Tv_{j}}+\sum_{j\in J}\overline{\left\langle Tv_{j},v\right\rangle}v_{j}\\ &=Tv-\sum_{j\in J}\left\langle Tv_{j},Tv\right\rangle Tv_{j}-\sum_{j\in J}\left\langle v_{j},Tv\right\rangle v_{j}\end{split}$
@@
Given a map $f:\mathrm{U}(1)\to\mathrm{U}(1)$, the winding number, denoted $[f]$, is an integer number exhibited by the natural isomorphism $\pi_{1}(\mathrm{U}(1))\simeq\mathbb{Z}$. It can be computed by considering the natural covering $g:\mathbb{R}\to\mathrm{U}(1)$ such that $g(t)=e^{i2\pi t}$. Then, the map $f$ can be lifted to a continuous map $\mu:[-\pi,\pi]\to\mathbb{R}$ such that $g(\mu(t))=f(\pi(t))$. Despite the numerous lifting options, the quantity $f]=\mu(\pi)-\mu(-\pi)\in\mathbb{Z}$ is well defined and depends only on the homotopy class of $f$.
@@
If $\lambda:\mathbb{T}^{1}\to\mathrm{U}(1)$ is a continuous map with $\lambda(\pi)=\lambda(-\pi)=\lambda(0)=1$ and such that $\lambda(-k)=\overline{\lambda(k)}$, then $\lambda]\in 2\mathbb{Z}$ is even.
@@
$\lambda]=\frac{\mu(\pi)-\mu(-\pi)}{2\pi}=\frac{2\mu(\pi)}{2\pi}\in 2\mathbb{Z}.\qed$
@@
In fact, the two branches of the function connect continuously for $k\to(t-1)\pi$, $k\to(1-t)\pi$ and also for $k\to\pi$, $k\to-\pi$ since $\lambda(\pi)=\lambda(-\pi)$. This means that $\lambda_{t}(k)$ is continuous. ∎
Proposition A.5.
Consider $\mathrm{U}(1)=\left\{e^{i\theta}\in\mathbb{C}\right\}$ as a topological group with the standard product of $\mathbb{C}$. If $f,g:\mathrm{U}(1)\to\mathrm{U}(1)$ are two smooth maps with $f(1)=g(1)=1$ then the winding number of $fg$ is the sum of the winding number of $f$ and the one of $g$.
@@
By [hatcher2005algebraic, Theorem 1.7], the map
$\mathbb{Z}\ni n\longmapsto\big[\alpha\mapsto e^{i2\pi n\alpha}\big]\in\pi_{1}(\mathrm{U}(1),1)$
is a group isomorphism. Moreover, if $f_{0}\simeq f_{1}$ and $g_{0}\simeq g_{1}$ are homotopies of based loops in $\mathrm{U}(1)$, then the pointwise product $f_{0}g_{0}$ is homotopic to $f_{1}g_{1}$, since multiplication in $\mathrm{U}(1)$ is continuous. Therefore, the assignment sending a loop to its winding number is compatible with pointwise multiplication of loops.
Applying this observation to the representatives $f$ and $g$, one concludes that the homotopy class of $fg$ corresponds to the sum of the homotopy classes of $f$ and $g$, and hence that the winding number of $fg$ is the sum of the winding numbers of $f$ and $g$. ∎
Theorem A.6 (Homotopy groups of $\mathrm{U}(n)$ [husemoller1966fibre, Chapter 8, Section 12]).
If $\mathcal{H}$ is a complex Hilbert space with finite dimension, then $\pi_{0}(\mathcal{U}(\mathcal{H}))\simeq\{0\}$, $\pi_{1}(\mathcal{U}(\mathcal{H}))\simeq\mathbb{Z}$ using the homomorphism that computes the winding number of the determinant, and $\pi_{2}(\mathcal{U}(\mathcal{H}))\simeq\{0\}$.
Theorem A.7 (Kuiper’s theorem [kuiper1965homotopy]).